Shift Summary, Owl 7/27/99

Angela Bellavance (bellavan@rice.edu)
Tue, 27 Jul 1999 07:00:19 -0500 (CDT)

Shift Summary for owl shift, 27/Jul/99

On Shift:

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of E832 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 7 (not written out)

Accounting of lost time: none

========================================================================
Runs ended this shift:
Run 14035 - COSMICS (not written out)
Run 14036 - MUON (crapped out, not useful at all)
Run 14037 - MUON (ongoing: 140 spills and nothing yet!...)

Tapes brought to FCC this shift: none
(including tapes for online split)

========================================================================
Accesses during shift: none
========================================================================
Overview of shift: waiting for beam, and waiting, and waiting, and...
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14035 00:00 Angela takes over duty of waiting for beam.
PeterSh hangs around as Leo hasn't shown up yet.
(This makes it twice in a row - is it me?
<sniffing armpits>...)
00:23 Message: "MI PS problems. No est. at this time.
Next update in one hour."
00:50 Leo arrives after being called out of bed. He
claims he thought he was on day shift.
03:03 MCR calls - get ready for beam! End cosmics run.
14036 1 03:16 Start muon run, but permit won't come back.
KFRENDs, power cycling the repeater, and resetting
crate 53 don't help. Call run manager, who
suggests scrapping the run and re-initing.
03:38 Message: "Preparing to send beam to users."
14037 1 03:45 Start muon run. Re-initing seems to have worked,
no beam yet.
05:41 Message: "MI to Tevatron transfer problems.
Experts enroute. No estimate."
06:15 Leo does a web search on "dumb stuff" and gets the
Microsoft homepage as a hit.
07:00 Hand-off to day shift.
========================================================================
Follow up Notes on Previous Shift

-haven't done muon run - no beam yet
-haven't turned magnet on to pos. - no muon run yet

========================================================================
Items to Follow up on Next Shift

-Elliott asks that the NCD terminals not be rebooted.
-Next shift needs to ask Leonard to order more drinking water.

_______
Angela Bellavance work: (630) 840-4144
bellavan@rice.edu home: (630) 548-3412
http://www.owlnet.rice.edu/~bellavan/ page: (630) 722-6861
"It has yet to be proven that intelligence has any survival value."
- Arthur C. Clarke